OpenStudy (anonymous):

THE RATIO OF 6 TO Y IS 2 TO 3 FIND Y

OpenStudy (gorv):

6/y = 2/3 \[6/y=2/3\]

OpenStudy (gorv):

multiply 3 both sides \[6*3/y=2*3/3\]

OpenStudy (anonymous):

6/y =2/3 this is a 3 times ration multipl 2/3 by 3/3 and you get your answer

OpenStudy (anonymous):

which is 9

OpenStudy (gorv):

18/y=2 y=18/2=9 y=9

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

6/Y=2/3 3/Y=1/3 Y=9